Given the following recurrence relation (M).


an = −4an−1 + 5an−2, a0 = 2, a1 = 8


The solution of (M) is:

a. an = 3 − (−5)

n

b. an = 3 + (5)

n

c. an = (3)

n − 5

d. None of these


1
Expert's answer
2021-08-03T13:10:42-0400

Characteristic equation:

"{k^2} + 4k - 5 = 0"

"D = 16 + 20 = 36"

"{k_1} = \frac{{ - 4 - 6}}{2} = - 5"

"{k_2} = \frac{{ - 4 + 6}}{2} = 1"

Then

"{a_n} = {C_1} \cdot {\left( { - 5} \right)^n} + {C_2} \cdot {1^n} = {C_1} \cdot {\left( { - 5} \right)^n} + {C_2}"

"{a_0} = 2,\,{a_1} = 8 \Rightarrow \left\{ {\begin{matrix}
{{C_1} + {C_2} = 2}\\
{ - 5{C_1} + {C_2} = 8}
\end{matrix}} \right. \Rightarrow {C_1} = - 1,\,{C_2} = 3"

Then

"{a_n} = - 1 \cdot {\left( { - 5} \right)^n} + 3 = 3 - {\left( { - 5} \right)^n}"

Answer:a. "{a_n} = 3 - {\left( { - 5} \right)^n}"